# Evaluating Prediction-based Interventions

### v1

@@ D.2 Experiment 1: Treatment Exposure Effect
We conduct an empirical study similar to that described in Section  5. Namely, for each study, we consider two different treatment assignments (to $n$ total units) that have the same treated and untreated units, but assign these treated units to different decision makers (and we assume each case is only seen by one judge in each treatment assignment). We consider $k$ judges, who are assumed to be identical, apart from their assigned cases (i.e. each judge possesses a uniform baseline bias of $b_{k}$). Specifically, both of them have decision making model as described in equation (2), where $f(z)=az$ is a linear function and $b_{1}=b_{2}=b$.

### v2

@@ D.2 Experiment 1: Treatment Exposure Effect
We conduct an empirical study similar to that described in Section  5. Namely, for each study, we consider two different treatment assignments (to $n$ total decision subjects) that have the same treated and untreated decision subjects, but assign these treated decision subjects to different decision makers (and we assume each case is only seen by one judge in each treatment assignment). We consider $k$ judges, who are assumed to be identical, apart from their assigned cases (i.e. each judge possesses a uniform baseline bias of $b_{k}$). Specifically, both of them have decision making model as described in equation (2), where $f(z)=az$ is a linear function and $b_{1}=b_{2}=b$.